OpenStudy (trantom):

Why is the graph of a cube so infinity-ridden? I have been trying it out on a 3D grapher and it seems to be "∞=x^∞+y^∞+z^∞

OpenStudy (trantom):

all of the question mark boxes are a glitch, its supposed to be \[\infty=x ^{\infty} + y ^{\infty} + z ^{\infty}\]

OpenStudy (trantom):

Obviously a computer can't deal with infinity, so I was using the 4000th power, and I noticed if I tried 1=x^infty+y^infty+z^infty it wasn't quite a cube, and if I tried 2 it was closer, and 3 gave the appearance of a cube I'm assuming thats supposed to be infinity

OpenStudy (perl):

computers have limits on numbers they use